# Holographic analysis of near-conformal dynamics and light dilaton

### v1

@@ 3 Toy model
  • We treat the field $X$ in the probe limit, assuming only quadratic action also in the IR, and set a generic IR boundary conditions:
@@
  • For the UV completion, we glue together directly solutions of $X$ with $\Delta_{\mathrm{IR}}=2+i\nu$ and $\Delta_{\mathrm{UV}}=3$, requiring continuity of the function and its first derivative. It means
    $\ell^{2}m_{X}(r)^{2}=\left\{\begin{array}[]{clc}\Delta_{\mathrm{UV}}(4-\Delta_{\mathrm{UV}})&=3&\qquad(r<r_{\mathrm{UV}})\\ \Delta_{\mathrm{IR}}(4-\Delta_{\mathrm{IR}})&=4+\nu^{2}&\qquad(r_{\mathrm{UV}}<r<r_{\rm IR})\end{array}\right.$ (34)
@@ 3.1 Scale hierarchy
As for the UV boundary condition at $r=r_{\rm UV}$, we can directly match the IR solution (36) with the UV solution (35). That is, we glue these solutions together at $r=r_{\mathrm{UV}}$, requiring the continuity of the function and its first derivative. It determines the phase $\alpha$ and the field value $X_{0}$ around the UV scale in terms of $m_{q}$ and $\sigma$ as
@@
where $n$ is an integer and we assume $2+A/B\gg\nu$ for the second line. Since $r_{\mathrm{IR}}>r_{\mathrm{UV}}$, if the quantity in the parenthesis is negative, it allows us to have a solution with $n=0$, yielding a non-Miransky scaling. If that quantity is positive, $n$ has to be equal to or greater than 1. For $n=1$, we may have approximately Miransky scaling, i.e. $\log(r_{\mathrm{IR}}/r_{\mathrm{UV}})\approx\pi/\nu$, assuming that $\alpha\sim\nu$ as (37) suggests. On the other hand, $n>1$ implies that the solution $X(r)$ has a node which is likely to yield an instability of the solution upon perturbation. We will discuss this in more detail below.
The above argument shows that generally, both Miransky scaling and non-Miransky scaling are possible depending on the IR physics parameterized by the parameter $A/B$. Let us look at more details separately for the cases with $m_{q}=0$ and $m_{q}\neq 0$ in the following subsections.
@@ 3.1.1 Massless quarks
However, we can also take $\nu$ to be finite but small, instead of taking $\nu\to 0$ directly. Then close to the critical values of $A/B$, where the scaling changes, special solution can be found. In particular, the case where $A+2B=0$ exactly is special. Then Eq. (39) tells us
@@ 3.1.2 Massive quarks
Using Eq. (47), Eq. (48), and Eq. (50), we can plot all possible values of the ratios $\hat{m}_{q}$ and $\hat{\sigma}$ as the scale hierarchy $r_{\mathrm{IR}}/r_{\mathrm{UV}}$ varies. This will lead to a spiral in the ($\hat{m}_{q},\hat{\sigma}$) plane which has been coined the Efimov spiral Efimov:1970zz ; Iqbal:2011aj . The spirals are shown for various choices of $A/B$ with $\nu=0.15$ in Figs. 1 and 2. Note that such spirals are common in holographic models (see, e.g., Iqbal:2011in ; Iqbal:2011aj ). In a typical setting (similarly as in this work), the spirals are driven by the violation of the BF bound at an fixed point. In models similar to the current setup, spirals have been considered in Jarvinen:2015ofa ; Arean:2016hcs ; BitaghsirFadafan:2018efw .
@@
Let us discuss the structure of the figures, starting from Fig. 1. We chose to express the ratio $A/B$ in terms of an angle, $A/B=\tan\phi$, and to vary the angle linearly. This choice does not set any of the boundary conditions in a special role. The colors vary from light blue (at $\phi=-1.52$, chosen to lie close to $\phi=-\pi/2$, i.e., Dirichlet boundary conditions) to dark blue (at $\phi=1.48$). Notice that for each value, there are actually two spirals, related by reflection with respect to the origin. Let us also note that
@@
The backgrounds with zero mass are found at the intersection points with the solid red lines. For each curve, i.e., each IR boundary condition, there are infinitely many such solutions, but only the first and in some cases second of the solutions is visible in the plotted area. We note that almost all such solutions show Miransky scaling with $\log\hat{\sigma}\approx-2\pi/\nu\approx-41.9$ (the light green band in the plot). However, one of the solutions shows essentially no scaling, and another shows nonstandard scaling. This solution with non-standard scaling is the finite-$\nu$ counterpart of the critical boundary conditions with $A/B=-2$ – actually for this curve we have $A/B\approx-2.066$.
@@
Let us then comment on the behavior to small but finite quark mass. Examples of the curves with constant quark mass are given as the dashed, dotted, and dot-dashed curves in Fig. 2. Notice that because signs of fermion masses are not observables, we include solutions with both positive and negative $\hat{m}_{q}$ in the curves. When the quark mass is tiny (e.g., $\hat{m}_{q}\ll\exp(-2\pi/\nu)$ for Miransky scaling solutions), it can be treated as small perturbation and the ratio of the UV and IR scales is essentially unaffected. An example of such a case is the red dashed curve in Fig. 2: with this mass value the regular solutions are found at the intersection points of the red dotted curve with the blue curves. When $\hat{m}_{q}$ grows and starts to be of the same order with $\hat{\sigma}$ (e.g., $\hat{m}_{q}\sim\exp(-2\pi/\nu)$ for Miransky scaling solutions), it can be no longer treated as a small perturbation and the ratio of the UV and IR scales is affected nontrivially (but Miransky scaling still holds) for those IR boundary conditions which show Miransky scaling at zero mass. An example of such a case is the red dot-dashed curve in Fig. 2. For even larger masses i.e., $\exp(-2\pi/\nu)\ll\hat{m}_{q}\ll 1$, the ratio between the UV and IR scales is determined by the quark mass, except for the few special cases which have nonstandard scaling behavior. And example of such is shown as the red dotted curve. For the backgrounds with standard Miransky scaling, these results agree with the analysis of Jarvinen:2015ofa .
@@
From Fig. 2 we also see that the solutions with non-Miransky scaling, in the parts of the curves that drop steeply, typically satisfy $\arctan(\hat{\sigma}/\hat{m}_{q})\approx\pi/4$ (or $-3\pi/4$) meaning that $\hat{m}_{q}\approx\hat{\sigma}$. From Eq. (37) we see that the UV phase $\alpha$ is large in this region, i.e., it deviates from the expected generic behavior $\alpha\sim\nu$. While these steep steps in the curves are expected to be an universal features that is present also in more complicated models than the simple toy model, the location of the drops is model dependent. It is also possible, that the drop happens near $\arctan(\hat{\sigma}/\hat{m}_{q})\approx\pi/2$ in some models with UV structure that differs from the toy model, which would mean that the nonstandard scaling solutions are found even at small quark masses for typical IR boundary conditions.
@@
where $\hat{F}$ was normalized such that $\delta\hat{F}=-\hat{\sigma}\delta\hat{m}_{q}$. From the first expression we see that $\hat{F}$ is nonmonotonic along the spiral: the expression even changes sign as the ratio $r_{\mathrm{UV}}/r_{\mathrm{IR}}$ varies. This is expected because at some point $\delta\hat{m}_{q}$ must change sign on any spiral. However since $\nu$ is small, the value of $\hat{F}$ is typically negative. Moreover, since the overall scaling factor $r_{\mathrm{UV}}^{4}/r_{\mathrm{IR}}^{4}$ dominates the parameter dependence along the spiral, the lowest free energies are generally found for solution with large $r_{\mathrm{UV}}/r_{\mathrm{IR}}$, i.e., least amount of scaling. This is seen clearly for the zero mass solutions from the second expression (56): the first term vanishes and the second term is monotonic in the ratio.
@@ 4 UV completion with flowing coupling constant
Let us then study a more general case in which, instead of the choice $V(\phi)=\Lambda_{5}$ of the toy model, consider a more general dilaton potential $V(\phi)$. The potential then gives rise to solutions of the dilaton, which depend on the holographic coordinate $r$. This “holographic RG flow” of the dilaton correspond to running of the coupling on the field theory side. The inclusion of the dilaton replaces the direct gluing of the near-boundary solutions ($r<r_{\mathrm{UV}}$) to the solutions in the walking regime ($r_{\mathrm{UV}}<r<r_{\mathrm{IR}}$), which we did in the toy model of the previous section, by a smooth but non-trivial flow.
In our setup, where we want to study the physics near the conformal transition to QCD, a complete theory is expected to include both an UV and an IR fixed points, which are realized through extrema of the dilaton potential. A possible choice for the UV fixed point (chosen in the IHQCD and V-QCD models Gursoy:2007cb ; Gursoy:2007er ; Jarvinen:2011qe ) is such that the UV fixed point is at $\phi=-\infty$ and the logarithmic UV running of the dilaton is driven by corrections to the potential that are suppressed by powers of $e^{\phi}$. A simpler choice is to set the UV fixed point at a finite value, so that the mass of the dilaton around the fixed point controls the dimension $\Delta_{g}$ of the dual operator $\mathrm{Tr}F^{2}$ (see, e.g., Gubser:2008ny ). For stable solutions with positive mass, the dimension satisfies $42<\Delta_{g}<$ 4. Note that the case $\Delta_{g}=4$ is physically motivated as it is the weak-coupling value of the dimension of the operator – this choice is obtained in the IHQCD model or by considering higher-order perturbations of the potential around the fixed point instead of a mass term (see Appendix F). For our purposes the precise value of $\Delta_{g}$ is, however, not important, as the physics we are interested in here mostly arises from the behavior of the solutions near the IR fixed point. Therefore we mostly consider the simple case with fixed point at a finite value of $\phi$ and nonzero mass, keeping $\Delta_{g}$ as a free parameter. Let us denote the fixed point values of the dilaton as $\phi_{\mathrm{UV}}$ and $\phi_{\mathrm{IR}}$.
We will first define the model explicitly. We then discuss how this more general setup is related to the toy model in the previous section. As we shall show, the simple UV completion of the toy model works as a characteristic example of the general case. For example, the phases of the tachyon oscillation, and how they depend on $\nu$, is qualitatively similar in the toy model as for general UV completions with nontrivial dilaton flow.
A complete model with proper gravity sector however, allows us to extend the analysis in other directions. In particular, we can study the flavor contributions to the one and two point functions of the Yang-Mills sector, and carry out a more precise analysis of the presence of light scalar states.
@@ 4.1 A concrete model with gravity and matter sectors
For simplicity, we will assume below that $\Delta_{\mathrm{UV}}$ takes the free field value for the $\bar{\psi}\psi$ operator, $\Delta_{\mathrm{UV}}=3$. We are mostly interested in the (walking) case where the IR dimension slightly violates the BF bound, i.e., $\Delta_{\mathrm{IR}}=2+i\nu$ with $\nu\ll 1$.
@@ 4.2 Solutions with scale separation and linking to the toy model
We will now argue that the results obtained by the simple toy model of Sec. 3 continue to hold at qualitative level for the more general action of Eqs. (72) and (73). We first discuss the results for the background and the Efimov spiral and then the fluctuations. Some assumptions are however needed: the general discussion below will only be valid in the presence of scaling, i.e., we must have $r_{\mathrm{UV}}\ll r_{\mathrm{IR}}$ in the limit $\nu\to 0$. Recall that in Sec. 3 we found that depending on the choice of model in the IR, the separation of UV and IR scales may also be absent. In such cases we cannot derive model independent results by using the techniques presented here.
@@ 4.2.1 Scales of the flow and the geometry
The two parts of the background solution may be formally obtained from a generic background solution at finite positive $\nu$ by taking the limit $\nu\to 0$ keeping either $r_{\mathrm{UV}}$ or $r_{\mathrm{IR}}$ fixed (see Sec. 5.3 and Appendix I in Arean:2013tja ). Near both the fixed points, the geometry is close to AdS5:
@@
In order to make these statements precise, let us then give explicit definitions for the UV and IR scales $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$. First, $r_{\mathrm{UV}}$ corresponds to UV scale around which $\phi$ starts to deviate from $\phi_{\mathrm{UV}}$ significantly, while around $r_{\mathrm{IR}}$ the field $X$ grows large so that its backreaction and the nonlinearities in the action for $X$ will become important. To make this explicit, we consider potential of the form
@@
where $40<\Delta_{g}<$ 4. Since the perturbative dimension of the dual operator, $F^{2}$ is four, it is natural to take $\Delta_{g}$ to be close to this value. The boundary expansion for the dilaton becomes

### v2

@@ 3 Toy model
  • We treat the field $X$ in the probe limit, assuming only quadratic action also in the IR, and set generic IR boundary conditions:
@@
  • For the UV completion, we glue together directly solutions of $X$ with $\Delta_{\mathrm{IR}}=2\pm i\nu$ and $\Delta_{\mathrm{UV}}=3$, requiring continuity of the function and its first derivative. It means
    $-\ell^{2}m_{X}(r)^{2}=\left\{\begin{array}[]{clc}\Delta_{\mathrm{UV}}(4-\Delta_{\mathrm{UV}})&=3&\qquad(r<r_{\mathrm{UV}})\\ \Delta_{\mathrm{IR}}(4-\Delta_{\mathrm{IR}})&=4+\nu^{2}&\qquad(r_{\mathrm{UV}}<r<r_{\rm IR})\end{array}\right.$ (36)
@@ 3.1 Scale hierarchy
As for the UV boundary condition at $r=r_{\rm UV}$, we can directly match the IR solution (38) with the UV solution (37). That is, we glue these solutions together at $r=r_{\mathrm{UV}}$, requiring the continuity of the function and its first derivative. It determines the phase, $\alpha$, and the field value, $X_{0}$, around the UV scale in terms of $m_{q}$ and $\sigma$ as
@@
where $n$ is an integer and we assume $2+A/B\gg\nu$ for the second line. Since $r_{\mathrm{IR}}>r_{\mathrm{UV}}$, if the quantity in the parenthesis is negative, it allows us to have a solution with $n=0$, yielding a non-Miransky scaling. If that quantity is positive, $n$ has to be equal to or greater than 1. For $n=1$, we may have approximately Miransky scaling, i.e. $\log(r_{\mathrm{IR}}/r_{\mathrm{UV}})\approx\pi/\nu$, assuming that $\alpha\sim\nu$ as (39) suggests. On the other hand, $n>1$ implies that the solution $X(r)$ has a node where the background field $X$ is vanishing. Such a vanishing background is likely to yield an instability of the solution upon perturbation so that the system eventually goes to the $n=1$ solution. We will show this more precisely later by computing free energy of each configuration. Note that the presence of a tower of unstable condensed solutions is not specific to our setup but a rather general consequence of BF bound violation (see, e.g., Kutasov:2011fr ; Iqbal:2011in ).
The above argument shows that, generally, both Miransky scaling and non-Miransky scaling are possible depending on the IR physics parameterized by the parameter $A/B$. Let us look at more details separately for the cases with $m_{q}=0$ and $m_{q}\neq 0$ in the following subsections.
@@ 3.1.1 Massless quarks
However, we can also take $\nu$ to be finite but small, instead of taking $\nu\to 0$ directly. Then close to the critical values of $A/B$, where the scaling changes, a special solution can be found. In particular, the case where $A+2B=0$ exactly is special. Then Eq. (41) tells us
@@ 3.1.2 Massive quarks
Using Eq. (49), Eq. (50), and Eq. (52), we can plot all possible values of the ratios $\hat{m}_{q}$ and $\hat{\sigma}$ as the scale hierarchy $r_{\mathrm{IR}}/r_{\mathrm{UV}}$ varies. This will lead to a spiral in the ($\hat{m}_{q},\hat{\sigma}$) plane which has been coined the Efimov spiral Efimov:1970zz ; Iqbal:2011aj . The spirals are shown for various choices of $A/B$ with $\nu=0.15$ in Figs. 1 and 2. Note that such spirals are common in holographic models (see, e.g., Iqbal:2011in ; Iqbal:2011aj ). In a typical setting (similarly as in this work), the spirals are driven by the violation of the BF bound at a fixed point. In models similar to the current setup, spirals have been considered in Jarvinen:2015ofa ; Arean:2016hcs ; BitaghsirFadafan:2018efw .
@@
Let us discuss the structure of the figures, starting from Fig. 1. We chose to express the ratio $A/B$ in terms of an angle, $A/B=\tan\phi$, and to vary the angle linearly. This choice does not set any of the boundary conditions to a special role. The colors vary from light blue (at $\phi=-1.52$, chosen to lie close to $\phi=-\pi/2$, i.e., Dirichlet boundary conditions) to dark blue (at $\phi=1.48$). Notice that for each value, there are actually two spirals, related by reflection with respect to the origin. Let us also note that
@@
The backgrounds with zero mass are found at the intersection points with the solid red lines. For each curve, i.e., each IR boundary condition, there are infinitely many such solutions, but only the first and in some cases second of the solutions is visible in the plotted area. We note that almost all such solutions show Miransky scaling with $\log\hat{\sigma}\approx-2\pi/\nu\approx-41.9$ (the light green band in the plot). However, one of the solutions shows essentially no scaling, and another shows non-standard scaling. This solution with non-standard scaling is the finite-$\nu$ counterpart of the critical boundary conditions with $A/B=-2$ – actually for this curve we have $A/B\approx-2.066$.
@@
Let us then comment on the behavior of small but finite quark mass. Examples of the curves with constant quark mass are given as the dashed, dotted, and dot-dashed curves in Fig. 2. Notice that because signs of fermion masses are not observables, we include solutions with both positive and negative $\hat{m}_{q}$ in the curves. When the quark mass is tiny (e.g., $\hat{m}_{q}\ll\exp(-2\pi/\nu)$ for Miransky scaling solutions), it can be treated as small perturbation and the ratio of the UV and IR scales is essentially unaffected. An example of such a case is the red dashed curve in Fig. 2: with this mass value the regular solutions are found at the intersection points of the red dotted curve with the blue curves. When $\hat{m}_{q}$ grows and starts to be of the same order with $\hat{\sigma}$ (e.g., $\hat{m}_{q}\sim\exp(-2\pi/\nu)$ for Miransky scaling solutions), it can be no longer treated as a small perturbation and the ratio of the UV and IR scales is affected non-trivially (but Miransky scaling still holds) for those IR boundary conditions which show Miransky scaling at zero mass. An example of such a case is the red dot-dashed curve in Fig. 2. For even larger masses i.e., $\exp(-2\pi/\nu)\ll\hat{m}_{q}\ll 1$, the ratio between the UV and IR scales is determined by the quark mass, except for the few special cases which have non-standard scaling behavior. And example of such is shown as the red dotted curve. For the backgrounds with standard Miransky scaling, these results agree with the analysis of Jarvinen:2015ofa .
@@
From Fig. 2 we also see that the solutions with non-Miransky scaling, in the parts of the curves that drop steeply, typically satisfy $\arctan(\hat{\sigma}/\hat{m}_{q})\approx\pi/4$ (or $-3\pi/4$) meaning that $\hat{m}_{q}\approx\hat{\sigma}$. From Eq. (39) we see that the UV phase $\alpha$ is large in this region, i.e., it deviates from the expected generic behavior $\alpha\sim\nu$. While these steep steps in the curves are expected to be universal features that are present also in more complicated models than the simple toy model, the location of the drops is model dependent. It is also possible, that the drop happens near $\arctan(\hat{\sigma}/\hat{m}_{q})\approx\pi/2$ in some models with UV structure that differs from the toy model, which would mean that the non-standard scaling solutions are found even at small quark masses for typical IR boundary conditions.
@@
where $\hat{F}$ was normalized such that $\delta\hat{F}=-\hat{\sigma}\delta\hat{m}_{q}$. From the first expression we see that $\hat{F}$ is non-monotonic along the spiral: the expression even changes sign as the ratio $r_{\mathrm{UV}}/r_{\mathrm{IR}}$ varies. This is expected because at some point $\delta\hat{m}_{q}$ must change sign on any spiral. However since $\nu$ is small, the value of $\hat{F}$ is typically negative. Moreover, since the overall scaling factor $r_{\mathrm{UV}}^{4}/r_{\mathrm{IR}}^{4}$ dominates the parameter dependence along the spiral, the lowest free energies are generally found for solutions with large $r_{\mathrm{UV}}/r_{\mathrm{IR}}$, i.e., least amount of scaling. This is seen clearly for the zero mass solutions from the second expression (58): the first term vanishes and the second term is monotonic in the ratio.
@@ 4 UV completion with flowing coupling constant
Let us then study a more general case in which, instead of the choice $V(\phi)=\Lambda_{5}$ of the toy model, consider a more general dilaton potential $V(\phi)$. The potential then gives rise to solutions of the dilaton, which depend on the holographic coordinate $r$. This “holographic RG flow” of the dilaton corresponds to running of the coupling on the field theory side. The inclusion of the dilaton replaces the direct gluing of the near-boundary solutions ($r<r_{\mathrm{UV}}$) to the solutions in the walking regime ($r_{\mathrm{UV}}<r<r_{\mathrm{IR}}$), which we did in the toy model of the previous section, by a smooth but non-trivial flow.
In our setup, where we want to study the physics near the conformal transition to QCD, a complete theory is expected to include both an UV and an IR fixed points, which are realized through extrema of the dilaton potential. A possible choice for the UV fixed point (chosen in the IHQCD and V-QCD models Gursoy:2007cb ; Gursoy:2007er ; Jarvinen:2011qe ) is such that the UV fixed point is at $\phi=-\infty$ and the logarithmic UV running of the dilaton is driven by corrections to the potential that are suppressed by powers of $e^{\phi}$. A simpler choice is to set the UV fixed point at a finite value, so that the mass of the dilaton around the fixed point controls the dimension $\Delta_{g}$ of the dual operator $\mathrm{Tr}G^{2}$ (see, e.g., Gubser:2008ny ). For stable solutions with positive mass, the dimension satisfies $42<\Delta_{g}<$ 4. Note that the case $\Delta_{g}=4$ is physically motivated as it is the weak-coupling value of the dimension of the operator – this choice is obtained in the IHQCD model or by considering higher-order perturbations of the potential around the fixed point instead of a mass term (see Appendix F). For our purposes the precise value of $\Delta_{g}$ is, however, not important, as the physics we are interested in here mostly arises from the behavior of the solutions near the IR fixed point. Therefore we mostly consider the simple case with fixed point at a finite value of $\phi$ and non-zero mass, keeping $\Delta_{g}$ as a free parameter. Let us denote the fixed point values of the dilaton as $\phi_{\mathrm{UV}}$ and $\phi_{\mathrm{IR}}$.
We will first define the model explicitly. We then discuss how this more general setup is related to the toy model in the previous section. As we shall show, the simple UV completion of the toy model works as a characteristic example of the general case. For example, the phases of the tachyon oscillation, and how they depend on $\nu$, is qualitatively similar in the toy model as for general UV completions with non-trivial dilaton flow.
A complete model with proper gravity sector, however, allows us to extend the analysis in other directions. In particular, we can study the flavor contributions to the one- and two-point functions of the Yang-Mills sector, and carry out a more precise analysis of the presence of light scalar states.
@@ 4.1 A concrete model with gravity and matter sectors
For simplicity, we will assume below that $\Delta_{\mathrm{UV}}$ takes the free field value for the $\bar{q}q$ operator, $\Delta_{\mathrm{UV}}=3$. We are mostly interested in the (walking) case where the IR dimension slightly violates the BF bound, i.e., $\Delta_{\mathrm{IR}}=2+i\nu$ with $\nu\ll 1$.
@@ 4.2 Solutions with scale separation and linking to the toy model
We will now argue that the results obtained by the simple toy model of Sec. 3 continue to hold at the qualitative level for the more general action of Eqs. (74) and (75). We first discuss the results for the background and the Efimov spiral and then the fluctuations. Some assumptions are however needed: the general discussion below will only be valid in the presence of scaling, i.e., we must have $r_{\mathrm{UV}}\ll r_{\mathrm{IR}}$ in the limit $\nu\to 0$. Recall that in Sec. 3 we found that depending on the choice of model in the IR, the separation of UV and IR scales may also be absent. In such cases we cannot derive model independent results by using the techniques presented here.
@@ 4.2.1 Scales of the flow and the geometry
The two parts of the background solution may be formally obtained from a generic background solution at finite positive $\nu$ by taking the limit $\nu\to 0$ keeping either $r_{\mathrm{UV}}$ or $r_{\mathrm{IR}}$ fixed (see Sec. 5.3 and Appendix I in Arean:2013tja ). Near both fixed points, the geometry is close to AdS5:
@@
In order to make these statements precise, let us then give explicit definitions for the UV and IR scales $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$. First, $r_{\mathrm{UV}}$ corresponds to UV scale around which $\phi$ starts to deviate from $\phi_{\mathrm{UV}}$ significantly, while around $r_{\mathrm{IR}}$ the field $X$ grows large so that its backreaction and the non-linearities in the action for $X$ will become important. To make this explicit, we consider potential of the form
@@
where $40<\Delta_{g}<$ 4. Since the perturbative dimension of the dual operator, $G^{2}$ is four, it is natural to take $\Delta_{g}$ to be close to this value. The boundary expansion for the dilaton becomes